OpenStudy (anonymous):

What is the area of a sector with radius equal to 8 and measure of arc equal to 300°?

OpenStudy (anonymous):

r=8 area=64pi measure of arc equal to 300° x:64pi=300:360 x=(64pi*300)/360=(160 pi)/3=167.55

OpenStudy (anonymous):

use the formula A = 1.2 * r^2 * theta where theta is in radians = 1/2 * 64 * 300pi/180 = 167.55